# A Reliable Framework for Human-in-the-Loop Anomaly Detection in Time Series

### v2

@@ Abstract
Time series anomaly detection is a critical machine learning task for numerous applications, such as finance, healthcare, and industrial systems. However, even high-performed models may exhibit potential issues such as biases, leading to unreliable outcomes and misplaced confidence. While model explanation techniques, particularly visual explanations, offer valuable insights to detect such issues by elucidating model attributions of their decision, many limitations still exist—They are primarily instance-based and not scalable across dataset, and they provide one-directional information from the model to the human side, lacking a mechanism for users to address detected issues. To fulfill these gaps, we introduce HILAD, a novel framework designed to foster a dynamic and bidirectional collaboration between humans and AI for enhancing anomaly detection models in time series. Through our visual interface, HILAD empowers domain experts to detect, interpret, and correct unexpected model behaviors at scale. Our evaluation with two time series datasets and user studies demonstrates the effectiveness of HILAD in fostering a deeper human understanding, immediate corrective actions, and the reliability enhancement of models.
Index Terms:
Human-AI collaboration, time-series, anomaly detection, visual interface, XAI.
I Introduction
Time series anomaly detection is a crucial technique in machine learning (ML), with broad applications such as identifying irregular financial transactions [1], monitoring abnormal health indicators in patient records [2], and preempting equipment failures in manufacturing [3]. Despite its importance, the effectiveness of these models can be compromised by biases—where models make decisions based on incorrect input features, resulting in problematic conclusions. To address this, eXplainable AI (XAI) techniques, especially visual explanations, are employed to shed light on models’ decision-making processes. However, these methods often focus on specific instances without scaling or generalizing insights across different data groups. Moreover, they solely allow a one-way information flow from models to users, without enabling users to verify and act upon the derived insights. To address these challenges, we introduce HILAD, a human-in-the-loop framework to enhance the robustness and reliability of anomaly detection models. Through a carefully-designed interface, our framework empowers experts to effectively detect, interpret, and rectify model biases, thereby bolstering model performance and enhancing human trust. In the following, we delve deeper into different facets of our work.
I-A Time Series Anomaly Detection Models
ML techniques for time series anomaly detection can be broadly categorized into unsupervised and supervised approaches. When ground-truth anomaly labels are unknown, unsupervised learning techniques are introduced to approximate future data points [4, 5, 6], where the deviations between the predicted and true values serve as anomaly indicators. On the other hand, if labeled data are available, supervised models can be trained as classifiers to directly predict anomaly labels [7, 8, 9]. Such models can often achieve higher accuracy and be utilized in critical areas like fraud detection or medical diagnostics [10, 11]. However, supervised models are generally more susceptible to bias issues, causing significant risks [12]. Considering this, in our work, we focus on enhancing supervised models with a CNN architecture, aiming to detect, understand, and mitigate hidden biases in such models with a human-in-loop approach.
I-B XAI for Time Series Anomaly Detection
XAI methods are essential for enhancing transparency and building trust for time series anomaly detection models. Model agnostic approaches, such as LIME [13] and SHAP [14], attribute the model’s prediction of a specific instance to its features according to a surrogate model and game theory, respectively. Without the need to access models’ internal parameters, these methods are more generalizable for explaining various models, but fall short of specificity, tending to produce low-quality results, especially for deep models [15, 16, 17]. Contrastively, methods designed for specific model architectures can often provide more precise explanations with higher quality [18, 19, 20], despite compromising the flexibility to explain a broader range of models. For example, CAM and its variants [21, 22] are widely adopted in CNN explanations because of their capabilities of providing direct and intuitive explanations through heatmaps, highlighting which parts of the data mostly correspond to the model’s decision [23]. Specifically, CAM can assign a significance score to each time step of an input, indicating their temporal importance to a specific model prediction. Fig. 1 shows two time sequences with “anomaly” labels and a model makes “correct” predictions for both, where the actual periods that the anomaly happens are highlighted in blue and the model’s decision attributions provided by CAM are highlighted in green. In this example, it is clear that the blue and green regions overlap in (a) but not in (b), which means the model attributes anomaly prediction to normal time periods and generates seemingly correct results in (b). Example in (b) showcases the harmful model bias issue, and how CAM can detect such hidden issues.
@@ I-B XAI for Time Series Anomaly Detection
Figure 1: Two sequences from the NASA MSL dataset with the ground truth label as “anomaly”. The true time frames corresponding to anomalies are highlighted in blue, and model’s attributes provided by CAM are highlighted in green. While the model makes “anomaly” prediction for both, regions highlighted by two colors overlap in (a) but not in (b), showing the model leverages the wrong features as the anomaly indicator in (b).
I-C Reliability Considerations of Models
Biased models can cause significant risks in real-world applications [24]. For instance, if a model identifies normal time series as anomalies, it underscores a fundamental misunderstanding of the underlying anomalous features, leading to potential overlooking of real anomalies [25]. While visual explanations such as CAM can reveal such discrepancies, their occurrence raises critical questions — Are these discrepancies a common feature across the entire data corpus, or are they isolated incidents? Moreover, once such biases are identified, what are the most effective methods for addressing and mitigating them? The discussion in Sec. I-B points out the limitations of current XAI techniques in fully tackling these challenges. This scenario necessitates further research and development of methods that not only highlight but also rectify these biases. Enhancing the reliability of anomaly detection models through improved understanding and correction of biases is essential to foster trust and ensure the efficacy of these systems in real-world applications.
I-D Potential of Human-AI Collaboration
Handling model reliability issues in time series anomaly detection is fundamentally a human-centric task. The complexity of determining what constitutes an anomaly necessitates a deep understanding that goes beyond algorithmic detection, requiring expert judgment and contextual awareness that machines alone are not equipped to provide. Researchers have actively developed visual analytics (VA) systems that incorporate human knowledge into the loop, designing visualization techniques to assist in the analysis of complex datasets and models in time series anomaly detection. However, current methods mainly involve single-instance inspections and do not provide effective summarization of predominant issues, thus limiting their ability to enhance models at scale. Additionally, these methods often lack sufficient interpretability support and require extensive manual verification by experts. To overcome these limitations, our work with HILAD aims to advance the state of the art by enhancing model interpretability, enabling the summarization and highlighting of critical data insights, thereby scaling up the integration of human knowledge into the machine learning process. HILAD is designed to streamline human-AI collaboration, making it more efficient at improving the accuracy and reliability of anomaly detection models.
I-E Our Contribution
@@
HILAD is a framework with an interactive interface, designed to enhance the reliability of ML models in the context of classification-based anomaly detection models. Our framework leverages visual explanation techniques to detect model biases, integrated with attribution-aware clustering and visual information summary to support group-level model issue interpretation and validation. The interface of HILAD navigates domain experts through the model validation workflow, facilitating the identification and mitigation of potential model biases. Simultaneously, HILAD’s spuriousness estimation mechanism allows humans to contribute domain knowledge at scale, where they manually annotate a few problematic clusters and then verify the estimated spuriousness of others. Lastly, we present a method for enhancing the model by rectifying identified errors based on human feedback. This bidirectional interaction supported by our framework utilizes human expertise effectively with a feedback loop, allowing humans to enhance AI systems with minimal effort.
@@
To demonstrate the effectiveness of HILAD, we invite domain experts to conduct model enhancement studies with benchmark datasets, and design comparative analyses to systematically compare our approach with state-of-the-art baselines through quantitative performance evaluation. Our evaluation encompasses both objective metrics and subjective assessments and aims to investigate the following hypotheses:
@@
  1. HILAD, a framework that integrates human expertise with ML techniques to enhance the reliability of time-series anomaly detection models, allowing for systematically detecting, interpreting, and mitigating model biases.
@@
  3. A solution with humans in the loop for generating actionable insights and incorporating such insights in model enhancement, which can inspire future research on human-assisted trustworthy AI.
II The Design of HILAD
@@
Figure 2: The HILAD framework involves three main phases: “Behavior Summarization”, “Annotation and Propagation”, followed by “Model Enhancement”. With a time-series (TS) dataset and an anomaly detection model, we obtain model attributions with CAM. At the “Behavior Summarization” phase, we conduct attribution-aware clustering to obtain clusters with both similar TS data features and consistent model attributions, and we compute cluster summarization to aggregate each cluster’s pattern with a single visualization. At the “Annotation and Propagation” phase, users interact with our designed VA interface to identify and annotate model bias issues with the help of Spuriousness propagation. Lastly, the annotation and user-verified “Spuriousness score” are used at the “Model Enhancement” phase, which employs a model regularization method to mitigate detected model bias.
II-A HILAD Framework
HILAD accepts a time-series (TS) dataset with ground-truth anomaly labels and a trained ML model for anomaly classification as input; provides an interactive interface for domain experts to detect and understand model issues; and supports issue mitigation with an integrated model regularization technique. Fig. 2 illustrates the framework of HILAD. In our quest for a model behavior interpretation that is intuitive and has the potential to be summarized, we employ an XAI technique, CAM [21]. With individual time sequences and an ML classifier for anomaly detection, CAM can produce visual explanations highlighting model attributions of its decisions. In our “Behavior Summarization” phase, we compute the similarity metrics of the original TS data and the model attributions, respectively, to conduct attribution-aware clustering from the latent space, as detailed in Sec. II-C. To enhance the efficiency of identifying model’s biased attribution issues, we further summarize each cluster with an integrated visualization, enabling experts to verify the attribution correctness without inspecting individual sequences. Our following “Annotation and Propagation” phase involves a visual interface for user interactions, where users overview clusters’ summarized model attributions, validate the exhibition of model issues, and annotate verified issues with a few clicks. As long as at least one cluster with “spurious attention” has been annotated, our backend method propagates the user’s assessment across other clusters as the estimated “Spuriousness scores”, enabling a facilitated model validation. Detailed explanations of this process can be found in Sec. II-D. After identifying the incorrect model’s attention, our model regularization strategy at the “Model Enhancement” stage systematically addresses these inaccuracies, refining the model’s focus to improve its overall performance and reliability, as described in Sec. II-E.
Figure 3: HILAD applied to the anomaly detection of a univariate time series classifier trained on the MSL dataset [26]. A System Menu, enabling the selection of dataset, class, and visualization options (Scatter plot view or Detailed plot view). B Cluster Information window, showing the data clusters and cluster metrics. C Representation Space window, showing a visual overview and the relative positions of all data clusters. D Instance Information window, showing individual time series sequences with model attributions belonging to a selected data cluster.
II-B System Interface
@@
The visual interface of HILAD comprises four main components. Fig. 3 depicts our interface when validating an anomaly classifier on the Mars Science Laboratory (MSL) dataset.
@@
The System Menu (Fig. 3  A) provides options for configuration selections, including dataset, class, and visualization layout. Our primary objective is to analyze the model’s comprehension of anomalous patterns; therefore, we have configured the default class setting to represent anomaly-related instances, with ground-truth or prediction class as an anomaly. The Cluster Information (Fig. 3  B) presents an explainable behavior summarization for the current model in each cluster. Each cluster aggregates similar instances that garner analogous attention from the model. On the top of Cluster Information window, we introduce cluster metrics to facilitate navigation through the dataset. The available metrics, including Accuracy, Confidence, and Spuriousness — the latter derived via the Spuriousness Propagation method (Sec. II-E) contingent upon user annotation — are presented to enhance the rapidity of analysis. Once click/annotate one cluster in Cluster Information window, its corresponding position in 2D space will be highlighted in the Representation Space window (Fig. 3  C) and instances in this cluster will be detailed in the Instance Information window (Fig. 3  D). Two visualization layouts are available for Representation Space (Fig. 3  C): Scatter plot view or Detailed plot view, allowing for an overarching view of the similarity of clusters in 2D space or a more detailed inspection of each cluster. This function is also embedded in Representation Space to facilitate users to switch more flexibly. We also provide information on the correctness of the current model in terms of classification and attention.
@@
II-C Behavior Summarization
At the “Behavior Summarization” phase, we first apply CAM to obtain model attributions correlated to its decision for each individual sequence. Then, we measure the similarities between: (1) the time-series data, according to the Dynamic Time Warping (DTW) Metric, and (2) the model attribution masks, according to the cosine similarity, respectively. With this information, we perform attribution-aware clustering to obtain clusters with similar data features and model attributions. To allow for effective user investigation, we visually summarize each cluster to help users identify common patterns of each cluster at a simple glance. We describe the details in the following.
II-C1 Time-series Context Similarity Calculation
In the realm of time series analysis, assessing the similarity between sequences is a pivotal task. A prevalent technique to measure this similarity is the DTW algorithm, which strives to minimize the distance between two time-sequenced datasets by dynamically aligning and “warping” them in the time dimension. A smaller DTW distance indicates a higher similarity. Formally, for two multivariate time series $T_{Q}=\{{T_{q}}_{1},{T_{q}}_{2},\dots,{T_{q}}_{n}\}$ and $T_{P}=\{{T_{p}}_{1},{T_{p}}_{2},\dots,{T_{p}}_{m}\}$, ${T_{q}}_{i}\in\mathbb{R}^{d}$ and ${T_{p}}_{j}\in\mathbb{R}^{d}$ , the DTW distance $d_{DTW}(T_{Q},T_{P})$ can be computed using the following recursive formula:

### v4

@@ Abstract
Time series anomaly detection is a critical machine learning task for numerous applications, such as finance, healthcare, and industrial systems. However, even high-performing models may exhibit potential issues such as biases, leading to unreliable outcomes and misplaced confidence. While model explanation techniques, particularly visual explanations, offer valuable insights by elucidating model attributions of their decision, many limitations still exist—They are primarily instance-based and not scalable across the dataset, and they provide one-directional information from the model to the human side, lacking a mechanism for users to address detected issues. To fulfill these gaps, we introduce HILAD, a novel framework designed to foster a dynamic and bidirectional collaboration between humans and AI for enhancing anomaly detection models in time series. Through our visual interface, HILAD empowers domain experts to detect, interpret, and correct unexpected model behaviors at scale. Our evaluation through user studies with two models and three time series datasets demonstrates the effectiveness of HILAD, which fosters a deeper model understanding, immediate corrective actions, and model reliability enhancement.
1 Introduction
Time series anomaly detection is a crucial technique in machine learning (ML), with broad applications such as identifying irregular financial transactions Ahmed et al. (2016), monitoring abnormal health indicators in patient records Šabić et al. (2021), and preempting equipment failures in manufacturing Li et al. (2019). Despite its importance, the effectiveness of these models can be compromised by biases—where models make decisions based on incorrect input features, resulting in problematic conclusions.
To tackle this issue, explainable AI (XAI) techniques Dwivedi et al. (2023); Xuan et al. (2024a) have been introduced to expose the decision logic of anomaly detectors. Among them, visual explanation methods like Class Activation Mapping (CAM) Zhou et al. (2016); Boniol et al. (2023) are particularly effective in attributing importance scores to each time step in a sequence. CAM can assign a significance score to each input time step, indicating its contribution to a model prediction. For example, Fig. 1 illustrates two time sequences with both labeled as “anomaly”, and the model makes correct predictions. However, CAM reveals a critical distinction: in case (a), the model’s attention (green) aligns well with the actual anomalous segment (blue), while for (b), the attribution of the prediction (green) does not overlap with the ground-truth (blue). Although both predicted labels seem correct, only the first reflects proper reasoning. This discrepancy demonstrates a hidden model bias, where decisions appear valid but are based on faulty logic, and highlights the need for an explainable system that not only exposes the bias but also helps mitigate such failures.
However, most existing XAI methods are limited in two key aspects: they analyze individual instances without generalizing to systematic model behaviors, and they enable only one-way interpretation—from model to user—without allowing users to intervene or improve the model Barkouki et al. (2023). In complex real-world environments, such limitations can lead to brittle deployments, where spurious model logic remains undetected and uncorrected at scale.
To ensure robust and trustworthy anomaly detection, we argue that explanation alone is insufficient. Instead, effective systems must support human-in-the-loop workflows where users can not only validate model decisions but also provide corrective feedback that influences model behavior. This perspective is supported by recent progress in human-AI collaboration, which shows how visual analytics interfaces can empower users to audit, refine, and debug AI models Amershi et al. (2019); Xuan et al. (2025a, b).
@@ 1 Introduction
Figure 1: Two sequences from the NASA MSL dataset with the ground truth label as “anomaly”. The true time frames corresponding to anomalies are highlighted in blue, and model’s CAM attributes are highlighted in green. While the model makes “anomaly” prediction for both, regions highlighted by two colors overlap in (a) but not in (b), showing the model leverages wrong features as the anomaly indicator in (b).
@@
In this paper, we introduce HILAD, a human-in-the-loop framework with an interactive interface, designed to enhance the reliability of ML models in the context of time series anomaly classification. Our framework leverages visual explanation techniques to detect model biases, designing attribution-aware clustering and summarization to foster model issue interpretation and validation. The interface of HILAD navigates domain experts through the model validation workflow, facilitating the identification and mitigation of model biases. Our spuriousness propagation mechanism allows humans to contribute domain knowledge at scale, where they only need to manually annotate a few clusters, and the estimated spuriousness scores of others are provided automatically for their verification. Lastly, we present a method for enhancing the model by rectifying identified errors based on human feedback. Our bidirectional interaction utilizes human expertise effectively with a feedback loop, allowing humans to enhance AI systems with minimal effort.
@@
To demonstrate the effectiveness of HILAD, we invite domain experts to conduct model enhancement studies with benchmark datasets, and design comparative analyses to quantitatively compare our approach with representative baselines. Our evaluation encompasses both objective metrics and subjective assessments and aims to investigate the following hypotheses:
@@
  1. HILAD, a framework that integrates human expertise with ML techniques to enhance the reliability of time series anomaly detection models, allowing for systematically detecting, interpreting, and mitigating model biases.
@@
  3. A human-AI-collaborative solution for generating actionable insights for model validation and enhancement, which can inspire future research on human-assisted trustworthy AI.
2 Related Works
2.1 Time Series Anomaly Detection Models
ML techniques for time series anomaly detection can be broadly categorized into unsupervised and supervised approaches. When ground-truth anomaly labels are unknown, unsupervised learning techniques are introduced to approximate future data points Kieu et al. (2019); Malhotra et al. (2015); Taylor and Letham (2018); Tuli et al. (2022), where the deviations between the predicted and true values serve as anomaly indicators. On the other hand, if labeled data are available, supervised models can be trained as classifiers to directly predict anomaly labels Del Campo et al. (2021); Gallicchio et al. (2017); Zhao et al. (2017). Such models can achieve higher accuracy and be utilized in critical areas like fraud detection or medical diagnostics Nabrawi and Alanazi (2023); Meena and Roy (2022). However, supervised models are generally more susceptible to bias issues, causing significant risks in practice Uddin et al. (2019). In our work, we focus on enhancing supervised models that directly predict anomaly labels, where various architectures such as transformers and CNN Dosovitskiy et al. (2020); Liu et al. (2021a); Boniol et al. (2023); Liu et al. (2021b); Yang et al. (2021) are fully supported, and we aim to detect, understand, and mitigate hidden biases in such models with a human-in-loop approach.
2.2 Reliability Considerations of Models
Biased models can cause significant risks in real-world applications Jung et al. (2021); Wang et al. (2024). For instance, if a model identifies normal time series as anomaly indicators, it underscores a fundamental misunderstanding of the underlying anomalous features, leading to potential overlooking of real anomalies Goswami et al. (2022); Deng and Kong (2021). While visual explanations such as CAM can reveal such discrepancies, their occurrence raises critical questions Xuan et al. (2024b) — Are these discrepancies a common feature across the entire data corpus, or are they isolated incidents? Moreover, once such biases are identified, what are the most effective methods for addressing and mitigating them? The discussion in Sec. 1 points out the limitations of current XAI techniques in fully tackling these challenges. This scenario necessitates further research and development of methods that not only highlight but also rectify these biases. Enhancing the reliability of anomaly detection models through improved understanding and correction of biases is essential to foster trust and ensure the efficacy of these systems in real-world applications.
2.3 Human-AI Collaboration
The growing field of human-AI collaboration has demonstrated the importance of integrating human expertise into machine learning workflows, particularly in high-stakes applications such as medical imaging, finance, and industrial monitoring Amershi et al. (2019); Schaekermann et al. (2020); Xuan et al. (2022); Zhang et al. (2023); Yildirim et al. (2024); Xuan et al. (2024b); Deng et al. (2025). Prior research in interactive AI systems has explored how visual analytics (VA) platforms can support experts in understanding, validating, and improving models, with much of the focus on vision and natural language processing tasks Yang et al. (2022, 2023); Xuan et al. (2025a, b); Li et al. (2022a); He et al. (2021); Gou et al. (2020); Liu et al. (2012), as these tasks are generally more interpretable to humans compared to time series data. For example, VA systems designed in Xuan et al. (2025a); Li et al. (2022a) allow users to iteratively refine and probe deep learning models. Similar approaches in object detection and segmentation Gou et al. (2020); He et al. (2021); Yan et al. (2025) and medical AI Xuan et al. (2024b); Malik et al. (2025) have enabled human-guided model adjustments. These works underscore the potential of bidirectional human-AI collaboration, where users not only interpret model outputs but actively refine them through feedback loops. In time series anomaly detection, existing approaches Geiger et al. (2020); Alnegheimish et al. (2022); Liu et al. (2022); Feng et al. (2023) facilitate interactive anomaly investigation by displaying feature attributions and model decisions. While these methods improve interpretability, they often rely on instance-level explanations, limiting their scalability to larger datasets Ruta et al. (2019). As a result, users must perform labor-intensive manual inspections, which are impractical in complex industrial or financial environments where anomalies are highly dynamic and demand rapid adaptation.
To address these challenges, we introduce HILAD to bridge the gap between visual interpretation and direct model refinement. HILAD supports a bidirectional interaction paradigm, where users can detect, interpret, and correct model misbehaviors at scale. Inspired by prior works that leveraged 2D projections for visualizing time series data Guo et al. (2021); Fujiwara et al. (2020), we incorporate a projection step to facilitate human interaction. However, unlike directly projecting raw features, we construct an attribution-aware latent space using an aggregated distance matrix over data and model attributions. HILAD leverages this space as an intermediate structure, supporting scalable annotation propagation and downstream model enhancement, which advances human-AI collaboration beyond prior visual-inspection-based approaches. By summarizing dominant failure patterns and allowing human adjustments, our system advances the integration of human expertise into AI-driven anomaly detection, enhancing model reliability and interpretability in time series applications.
3 HILAD Framework
@@
Figure 2: The HILAD framework involves three phases: “Behavior Summarization”, “Annotation and Propagation”, and“Model Enhancement”. At the “Behavior Summarization” phase, we conduct attribution-aware clustering, where each cluster includes TS instances exhibiting similar data features and model attributions. We also visualize the overall pattern of each cluster to streamline user interpretation. At the “Annotation and Propagation” phase, users interact with our designed VA interface to identify and annotate model bias issues. These annotations are propagated to unannotated clusters to generate “Spuriousness scores” for user verification. Lastly, we mitigate the detected issues at the “Model Enhancement” phase.
HILAD is a human-in-the-loop framework designed to improve the reliability of anomaly detection models on time series (TS) data. It takes as input a trained anomaly detection model along with TS instances with ground-truth anomaly labels. The system then enables domain experts to interactively inspect, validate, and correct model behaviors through three coordinated phases: “Behavior Summarization”, “Annotation and Propagation”, and “Model Enhancement”, as shown in Fig. 2.
The core idea of HILAD is to assist users in efficiently detecting and addressing model misbehavior caused by spurious attributions, where the model relies on misleading or irrelevant signals to make decisions. To support this, in the first “Behavior Summarization” phase, HILAD automatically generates and visualizes clusters with similar data features and model attributions. As illustrated in Fig. 2, each cluster represents a group of TS sequences that share both contextual similarity in terms of data signals and model attribution similarity, and such patterns are visualized in “cluster summarization”, enabling experts to verify the attribution correctness without inspecting individual sequences. Then, in the following “Annotation and Propagation” phase, we enable users to overview clusters’ summarized model attributions, validate the exhibition of model issues, and annotate verified issues with a few clicks, where we provide label propagation to facilitate the validation process. Finally, the “Model Enhancement” stage systematically addresses these inaccuracies, refining the model’s focus to improve its overall performance and reliability.
3.1 System Interface
@@
The visual interface of HILAD comprises four main components. Fig. 3 depicts our interface when validating an anomaly classifier on the Mars Science Laboratory (MSL) dataset Hundman et al. (2018).
@@
The System Menu (Fig. 3  A) provides options for configuration selections, including dataset, class, and visualization layout. Our primary objective is to analyze the model’s comprehension of anomalous patterns; therefore, we have configured the default class setting to represent anomaly-related instances, with ground-truth or prediction class as an anomaly. The Cluster Information (Fig. 3  B) presents an explainable behavior summarization for the current model in each cluster. Each cluster aggregates similar instances that garner analogous attention from the model. On the top of Cluster Information window, we introduce cluster metrics to facilitate navigation through the dataset. The available metrics, including Accuracy, Confidence, and Spuriousness — the latter derived via the Spuriousness Propagation method (Sec. 3.4) contingent upon user annotation — are presented to enhance the rapidity of analysis. Once click/annotate one cluster in Cluster Information window, its corresponding position in 2D space will be highlighted in the Representation Space window (Fig. 3  C) and instances in this cluster will be detailed in the Instance Information window (Fig. 3  D). Two visualization layouts are available for Representation Space (Fig. 3  C): Scatter plot view or Detailed plot view, allowing for an overarching view of the similarity of clusters in 2D space or a more detailed inspection of each cluster. This function is also embedded in Representation Space to facilitate users to switch more flexibly. We also provide information on the correctness of the current model in terms of classification and attention.
Figure 3: HILAD applied to the anomaly detection of a univariate time series classifier trained on the MSL dataset Hundman et al. (2018). A System Menu, enabling the selection of dataset, class, and visualization options (Scatter plot view or Detailed plot view). B Cluster Information window, showing the data clusters and cluster metrics. C Representation Space window, showing a visual overview and the relative positions of all data clusters. D Instance Information window, showing individual time series sequences with model attributions belonging to a selected data cluster.
@@
3.2 Behavior Summarization
At the “Behavior Summarization” phase, for each individual instance, we first apply CAM to obtain model attributions correlated to its decision, providing a one-dimensional vector (or mask) with the same length as the time sequence. Each value in this vector falls within the range (0, 1) and reflects the importance of the corresponding timestamp to the model’s prediction. Then, we measure and aggregate similarities between: (1) TS data, according to the Dynamic Time Warping (DTW) Metric, and (2) model attribution masks, according to the cosine similarity, respectively. With this information, we perform attribution-aware clustering to obtain clusters with similar data features and model attributions. To allow for effective user investigation, we visually summarize each cluster to help users identify common patterns of each cluster at a simple glance.
Note that CAM is fully modular and can be replaced with a variety of attribution methods, including Grad-CAM Selvaraju et al. (2020), ScoreCAM Wang et al. (2020), and model-agnostic techniques like RISE Petsiuk et al. (2018). Our design does not rely on CAM-specific internals, allowing users to plug in alternative attribution mechanisms depending on the accessibility or specific requirements. We use CAM as our default XAI method, as CAM and CAM-derived techniques remain among the most widely used in time series and other domains due to their simplicity, interpretability, and computational efficiency. And they have been extensively validated across a broad range of models and datasets.
3.2.1 Time series Context Similarity Calculation
To measure the similarity between time sequences, we use the prevalent technique, the DTW algorithm, which strives to minimize the distance between two time-sequenced datasets by dynamically aligning and “warping" them in the time dimension. A smaller DTW distance indicates a higher similarity. Formally, for two multivariate time series $T_{Q}=\{{T_{q}}_{1},{T_{q}}_{2},\dots,{T_{q}}_{n}\}$ and $T_{P}=\{{T_{p}}_{1},{T_{p}}_{2},\dots,{T_{p}}_{m}\}$, ${T_{q}}_{i}\in\mathbb{R}^{d}$ and ${T_{p}}_{j}\in\mathbb{R}^{d}$ , the DTW distance $d_{DTW}(T_{Q},T_{P})$ can be computed using the following recursive formula: